Hi im having problems on this absolute value problems
|x - 2| + 10 = 12


Answer by stanbon(75887)

|x - 2| + 10 = 12
|x-2| = 2
x-2 = 2 or x-2=-2
x = 4 or x = 0
======================
Cheers,
Stan H.